OpenStudy (henrysteels):

What is an equation in point-slope form for the given point and slope? Point (2-6) ; Slope -3/4

OpenStudy (will.h):

Welcome to OpenStudy

OpenStudy (henrysteels):

Thanks!

OpenStudy (will.h):

Use the following form Y - y1 = m(x - x1) Where m is the slope Y1 = -6 and x1 = 2 Substitute and you'll have your answer

likeabossssssss (likeabossssssss):

welcome to os

OpenStudy (henrysteels):

is it y-6=-3/4(x-2)?

OpenStudy (henrysteels):

and thanks

OpenStudy (will.h):

No

OpenStudy (will.h):

Y - (-6) = -3/4(x - 2) Simplify the left side of the equation Y + 6 = -3/4(x -2)